3 students have 400$ altogetfer . How much does each have if the 2 students have altogether 286$ and the 2 other students have 230$ ?
Answer by oscargut(2103)

let a,b,c the amounts
then
1) a+b+c=400
2) a+b=286
3) b+c=230
using the first equation and the second 286+c=400 then c=114
using the third equation b+114=230 then b=116 then a+116=286
then a=170
R: amounts are 170,116 and 114